OpenStudy (anonymous):

If a line with a slope of 4 crosses the y-axis at (0, 1), what is the equation of the line?

OpenStudy (freckles):

y=mx+b where m is slope and (0,b) is the y-intercept

OpenStudy (freckles):

you are given both m and b

OpenStudy (freckles):

just plug them into y=mx+b

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ % (a,b) &({\color{red}{ 0}}\quad ,&{\color{blue}{ 1}})\quad \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= 4 \\ \quad \\ % point-slope intercept y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (ineedhelpfast12):

y=4x+1